# Value-Based Large Language Model Agent Simulation for Mutual Evaluation of Trust and Interpersonal Closeness

### v1

@@ Methods
Furthermore, we establish a condition in which a definition of value is added to the prompt based on proposals from related studies to reduce the ambiguity of abstract concepts, such as values [25]. The values were defined based on previous studies [26]. For instance, when controlling for power, the second-person prompt with a definition is as follows: “You are a person attributing extremely more importance to power (social status and prestige, control or dominance over people and resources (social power, authority, wealth)).” Additionally, we compared two methods of prompt type: providing the prompt as a system prompt and including it at the beginning of the user prompt.
@@
For the experimental procedure, the 40 portrait items were independently presented to the LLMs in random order to generate responses. This process was repeated 50 times to ensure robustness. Considering that extracting subjective information from LLMs can be challenging [29], if valid responses were not obtained for 10% or more of the items in the first iteration, the trial for that model and prompt combination was aborted, and the score was not calculated.
@@
    Relationship between Value Similarity and Mutual Evaluation: We defined the value similarity for each agent pair based on Schwartz’s circular model of values [6, 7]. In this model, values close to the circular continuum are motivationally compatible, whereas distant values are in conflict. Based on this principle, we classified value similarity into the following three levels. We analyzed the impact of value similarity on evaluations by comparing mutual evaluation scores (trust and interpersonal closeness) across the three similarity levels.
@@
    Correlation Analysis between Basic Value and Higher-Order Value Results: We verified whether controlling for a higher-order value reflects the collective tendencies of its corresponding basic values. To achieve this, we first created an aggregated higher-order value evaluation matrix by averaging the evaluation scores of the basic value pairs belonging to each higher-order dimension using the results from the basic value control experiments. Next, we calculated Pearson’s correlation coefficient between this aggregated matrix and the direct higher-order value evaluation matrix obtained by directly controlling for the higher-order values. Each evaluation matrix was flattened into a one-dimensional vector to calculate the correlation.
@@
Data will be made available on reasonable request.
Acknowledgements
This work was supported by JSPS KAKENHI Grant Number 22K17949, JST PRESTO Grant Number JPMJPR23I2, and JST BOOST, Japan Grant Number JPMJBS2402.
Author contributions statement
All authors (Y.S., T.U., H.I.) contributed to the study’s conception and design. Y.S. collected data, and Y.S. and T.U. analyzed the results. Y.S. and T.U. prepare the material, and all authors approved the final manuscript.
Competing interests
The authors declare no competing interests.

### v2

@@ Methods
Furthermore, we establish a condition in which a definition of value is added to the prompt based on proposals from related studies to reduce the ambiguity of abstract concepts, such as values [35]. The values were defined based on previous studies [36]. For instance, when controlling for power, the second-person prompt with a definition is as follows: “You are a person attributing extremely more importance to power (social status and prestige, control or dominance over people and resources (social power, authority, wealth)).” In addition, when controlling for self-enhancement, the third-person prompt without definitions is as follows: “The following are answers from a person attributing extremely more importance to power, achievement.” Additionally, we compared two methods of prompt type: providing the prompt as a system prompt and including it at the beginning of the user prompt.
@@
For the experimental procedure, the 40 portrait items were independently presented to the LLMs in random order to generate responses. This process was repeated 50 times to ensure robustness. Each API call to the LLM is independent, and the conversation history variables are reset each time, so that no systematic drift could occur within the same foundational LLM. Considering that extracting subjective information from LLMs can be challenging [39], if valid responses were not obtained for 10% or more of the items in the first iteration, the trial for that model and prompt combination was aborted, and the score was not calculated.
@@
In each dialogue, the two agents were implemented as independent calls to LLM. For every turn, system prompt was set separately, and no session-level hidden states or memory were shared across agents. Instead, the dialogue history was explicitly reconstructed at each step and inserted into the prompt for the relevant agent. This design ensured that each agent’s utterance was generated independently, while still maintaining conversational coherence through the prompt. Furthermore, no memory or hidden states were shared across the two dialogue tasks.
@@
    Relationship between Value Similarity and Mutual Evaluation: It is known that Schwartz’s values are not independent but rather possess a circular structure of motivational relationships. Therefore, in this study, we did not treat each value as independent and instead defined the value similarity between agents based on this circular model [6, 7]. In this model, values close to the circular continuum are motivationally compatible, whereas distant values are in conflict. Based on this principle, we classified value similarity into the following three levels. We analyzed the impact of value similarity on evaluations by comparing mutual evaluation scores (trust and interpersonal closeness) across the three similarity levels.
@@
    Correlation Analysis between Basic Value and Higher-Order Value Results: We verified whether controlling for a higher-order value reflects the collective tendencies of its corresponding basic values. To achieve this, we first created an aggregated higher-order value evaluation matrix by averaging the evaluation scores of the basic value pairs belonging to each higher-order dimension using the results from the basic value control experiments. For example, the aggregated score for the “self-enhancement to self-transcendence” dimension was calculated by averaging the mutual evaluation scores from the four corresponding basic value pairs: power to universalism, power to benevolence, achievement to universalism, and achievement to benevolence. Next, we calculated Pearson’s correlation coefficient between this aggregated matrix and the direct higher-order value evaluation matrix obtained by directly controlling for the higher-order values. Each evaluation matrix was flattened into a one-dimensional vector to calculate the correlation.
@@
The experimental code, prompts, and raw data analyzed in this study are available from the following repository: https://osf.io/7jn8x/?view_only=a1afa8564ac04a4481ccb03b3f0dd555